Dipa has 7 more Rs.2 coins than 50 paisa coins. If the value of all the coins is Rs 76.50, find the number of each type of coins.
Answer by mananth(16949)

Dipa has 7 more Rs.2 coins than 50 paisa coins. If the value of all the coins is Rs 76.50, find the number of each type of coins.
..
let 50 paise coins she has be x
Rs.2 coins will be x+7
..
50 paise = Rs. 0.5
..
Value Equation
0.5x+2(x+7)=76.50
0.5x+2x+14 = 76.50
2.50x=76.50-14
2.50x=62.50
divide by 2.50
x=62.50/2.50
x= 25 coins of 50 paise
25+7 = 32 coins of Rs.2
..
CHECK
32*2+25*0.5=76.5
...
m.ananth@hotmail.ca